OpenStudy (anonymous):

6(r/r-1)^2 - 7(r/r-1) - 5 = 0. Find all real solutions ( let w = r/r-1)

OpenStudy (anonymous):

Is r= -5/3 and r = 1/3 correct?

OpenStudy (shamil98):

w = -1/2 , w = 5/3 sub back in to solve for r.. r/r-1 = -1/2 r = -1/2(r-1) r = -1/2r + 1/2 3/2 r = 1/2 r = 1/12 r/r-1 = 5/3 r = 5/3(r-1) r = 5/3 r - 5/3 -2/3 r = - 5/3 r = 5/2